the slope-intercept equation for the line that passes through (-8, -11) and is perpendicular to -2x – 7y = -14
Answer by nerdybill(7384)

the slope-intercept equation for the line that passes through (-8, -11) and is perpendicular to -2x – 7y = -14
.
Determine slope of:
-2x – 7y = -14
–7y = 2x-14
y = (-2/7)x + 2
slope is -2/7
.
A line perpendicular to above must have a slope that is a "negative reciprocal":
(-2/7)m = -1
m = 7/2
.
plug given point and slope into "point-slope" form:
y - y1 = m(x - x1)
y - (-11) = (7/2)(x - (-8))
y + 11 = (7/2)(x + 8)
y + 11 = (7/2)x + 28
y = (7/2)x + 17 (this is what they're looking for)